A man alternately tosses a coin and throws a die beginning with the coin. Find the probability he gets a head in the coin before he gets a 5 or 6 in the die. ANSWER-3/4

Grade:12

1 Answers

Sher Mohammad IIT Delhi
10 years ago
.5+.5*4/6*.5+5*4/6*.5 *4/6*.5.............
.5(1+2/6+ (2/6)^2..........)=.5* 1/(1-2/6)=3/4
sher mohammad
b.tech, iit delhi